OpenStudy (anonymous):

Which expressions would complete this equation so that it has one solution? 4(2x – 3) – 4x = Choose exactly two answers that are correct. A. 4x – 12 B. 3x – 4 C. 3(2x – 6) D. 4(x – 5)

OpenStudy (phi):

can you distribute the 4 in : 4(2x – 3) (that means multiply 4 times each term inside the parens)

OpenStudy (phi):

refresh your browser if that does not display

OpenStudy (anonymous):

ok, @phi

OpenStudy (anonymous):

@r_josh10

OpenStudy (phi):

can you distribute the 4 in 4(2x-3) ?

OpenStudy (anonymous):

ok

OpenStudy (r_josh10):

Its B and C. Let's solve your equation step-by-step. 4x−12=3(2x−6) Step 1: Simplify both sides of the equation. 4x−12=6x−18 Step 2: Subtract 6x from both sides. 4x−12−6x=6x−18−6x −2x−12=−18 Step 3: Add 12 to both sides. −2x−12+12=−18+12 −2x=−6 Step 4: Divide both sides by -2. −2x −2 = −6 −2 x=3 Answer: x=3 b. Let's solve your equation step-by-step. 4x−12=3x−4 Step 1: Subtract 3x from both sides. 4x−12−3x=3x−4−3x x−12=−4 Step 2: Add 12 to both sides. x−12+12=−4+12 x=8 Answer: x=8